# Exploiting Domain-Specific Parallel Data on Multilingual Language Models for Low-resource Language Translation

### v1

@@ Abstract
keywords:
Neural Machine Translation , Domain divergence , multilingual Language Models , domain adaptation[label1]organization=School of Mathematical and Computational Sciences,addressline=Massey University, city=Albany, postcode=0632, state=Auckland, country=New Zealand
1 Introduction
@@
The need to rapidly develop translation systems to disseminate information during pandemics, natural disasters, and other major events justifies the need to build domain-specific Neural Machine Translation (NMT) systems with any and all available parallel data. This is particularly true for low-resource languages (LRLs) that do not possess large amounts of parallel data for any single domain [1]. Therefore, when building a domain-specific NMT model for a given LRL pair, we face two possible scenarios: the considered (target, aka test) domain does not have any parallel data to be used in training, or it has a little amount of parallel data. In both cases, it is possible to make use of parallel data available for the same language pair, but in different (auxiliary) domains. This is not an unrealistic assumption - according to Tiedemann [2], hundreds of languages have parallel data from more than one domain.
@@
Multilingual sequence-to-sequence Language Models (msLMs) such as mBART [3] have been shown to outperform vanilla Transformer models for the LRL-NMT task [4, 5, 6]. There are two possible ways to exploit auxiliary domain parallel data in building domain-specific LRL-NMT systems on top of msLMs: to further pre-train the msLM with self-supervised training objectives (continuous pre-training) [7] or to fine-tune the msLM with the NMT objective.
@@
As shown in Figure 1, there are different ways to use auxiliary domain data in fine-tuning. These can be categorized as single-stage fine-tuning (FT) and Intermediate Task Transfer Learning (ITTL). Similarly, we can introduce another categorization based on the data availability of the target domain. In-domain refers to the case whether target domain has some parallel data for model training. Out-domain refers to the case where there is no such parallel data for the target domain. Note that out-domain testing scenario can be considered as zero-shot testing with respect to the target domain.
@@
Variations of single-stage FT are: 1) single domain (vanilla) FT, where the msLM is fine-tuned once with data from one domain, and 2) multi-domain FT, where the msLM is fine-tuned once with data from multiple domains. In vanilla FT, if the target domain has parallel data, this data is used for fine-tuning (in-domain case). In multi-domain FT, when the target domain has parallel data (in-domain case), it is combined with auxiliary domain data. In ITTL, the msLM is fine-tuned with an intermediate task, before fine-tuning with the final task. It also has two variations: single-domain ITTL, where the intermediate task has data from one auxiliary domain, and multi-domain ITTL, where the intermediate task has data from multiple domains. In both cases, the final task (i.e. stage111We use the term stage interchangeably with task, i.e. final task or final stage, and intermediate task or intermediate stage.) fine-tuning is done using data from a single domain. If target domain has parallel data (in-domain case), the final task always uses this data. In the in-domain case, target domain data is also combined with auxiliary domain data in intermediate stage of multi-domain ITTL.
@@
Currently, there is no comparative study on the impact of continuous pre-training and fine-tuning techniques when auxiliary domain data is used. Determining the method that best exploits auxiliary domain parallel data is beneficial for LRL-NMT researchers who are unable to run extensive experiments due to limited computational resources. Lee et al. [6] and Adelani et al. [8] identified that domain divergence in parallel datasets affects NMT performance. However, they did not conduct a detailed investigation into the impact of this factor. Khiu et al. [9] and our previous work [10] evaluated the impact of domain divergence only in the case of single domain ITTL. Given the above shortcomings, this paper intends to answer the following questions related to domain-specific LRL-NMT systems built on top of msLMs:
1.
@@
In order to answer these questions, we conducted an extensive set of experiments with the aforementioned FT techniques as well as the bitext denoising pre-training technique [11]. We considered the LRLs Sinhala (si), Tamil (ta), Kannada (ka), and Gujarati (gu) as well as Hindi (hi) for English-centric NMT. We report results for the case where the target domain has some parallel data for training (in-domain) and where the target domain has no parallel data (out-domain). We quantify the impact of domain divergence using Jenson-Shanon Divergence (JSD).
@@
When the parallel data set size is small (less than 50k, in our experiments), pre-training with bitext denoising yields no gains for both in- and out-domain setups.
2.
@@
Multi-domain ITTL is the best for in-domain, but this utility diminishes when target dataset size increases. For out-domain, the best strategy depends on the domain relatedness and the size of parallel corpora.
3.
@@
For both in- and out-domain setups, simply combining data from multiple domains does not help in ITTL, mainly due to the divergence between the domains they belong to.
2 Related Work
2.1 ITTL on msLMs for NMT
@@
ITTL has roots in Transfer Learning, which has been extensively explored for NMT with vanilla Transformer models. However, as mentioned earlier, for LRLs, NMT systems built on top of msLMs outperform these traditional NMT techniques. Thus we do not discuss these techniques.
@@
ITTL [12, 13, 14] refers to fine-tuning a pre-trained language model with the intermediate task(s), before fine-tuning on the final task. It has been extensively experimented with encoder-based LLMs such as BERT [12, 15], and has shown promising results even for LRLs [16].
@@
With respect to ITTL for NMT, the closest to our work is Adelani et al. [8]. Similar to us, they explored single-domain ITTL and multi-domain ITTL, as well as multi-domain FT. However, they only considered the scenario where the auxiliary task has a large amount of data, while the target task has only a small amount of data. Moreover, they only used religious and news data and tested only for the in-domain setup. They did not quantify the impact of domain divergence. Our prior work [10] and Khiu et al. [9] considered only single-domain ITTL. However, they considered both in- and out-domain test setups and investigated the impact of domain divergence. Verma et al. [17] experimented with single-domain ITTL and multilingual ITTL (data from the same domain but in multiple languages). They found that the former is better, but did not quantify the impact of domain divergence.
2.2 Pre-training msLMs with Parallel Data
@@
Reid and Artetxe [7] augmented the existing denoising objective in mBART with three new objectives: replace words in the noised sequence with a bilingual dictionary, predict the reference translation instead of the input sequence (bitext denoising), and a combination of the two. Kale et al. [18] used four denoising objective functions on mT5: translation language modeling, standard NMT, denoised NMT and denoised NMT with language model (LM). Chi et al. [19] used three cross-lingual objective functions on mT5: machine translation, translation pair span corruption, and translation span corruption. However, the latter two research did not consider the NMT task.
3 Methodology
@@
For continuous pre-training, we experimented with two of the Reid and Artetxe’s pre-training objective functions: bitext de-noising with auxiliary domain parallel data with and without monolingual denoising [11]. For monolingual denoising, we used the source and target side of the parallel data separately and did not use additional monolingual data as done by Reid and Artetxe [11], because our objective is only to exploit the use of parallel data. We also excluded their dictionary denoising, as many LRLs still do not have sufficient bilingual dictionaries [20]. Once msLM is pre-trained with Reid and Artetxe’s objective functions [11], we continue with vanilla FT with the NMT objective function.
@@
We use the four FT strategies shown in Figure 1, and we formally define them as follows: Let be the set of all domains where there is parallel data available for model training, for a considered language pair. is a domain for which there is no training data, but only test data is available. In the single-domain FT (vanilla FT) baseline (Figure 1(a)), the model is only fine-tuned with data from one domain . In the in-domain case, both fine-tuning and testing is done using data from domain . In the out-domain case, fine-tuning uses data from , but testing is done on domain . In multi-domain FT (Figure 1(b)), the model is fine-tuned with data from multiple domains . In the in-domain case, this includes data from the target domain as well. In the out-domain case, this trained model is tested on the unseen target domain . In single-domain ITTL (Figure 1(c)), msLM is first fine-tuned with data of domain . This is the intermediate task, which is also called the first stage. Next, this fine-tuned model is further fine-tuned with data from domain . This is the second and last stage222The model can always be fine-tuned in more stages if there is data from different domains. However, we do not consider this option due to lack of computational resources.. In the in-domain case, testing will be done on the same domain . In out-domain case, testing will be done on the unseen domain . In multi-domain ITTL (Figure 1(d)), msLM is first fine-tuned with data from multiple domains in . This model is further fine-tuned with , which is the target domain. In the in-domain case, data of is included in the intermediate stage, as well as in the final stage. Similar to previous setups, in the out-domain case, testing will be done on the unseen domain
4 Experimental Setting
Our experiment setup is inspired from our previous work [10] and we follow the same choices for msLM, languages, datasets, etc.
msLM: We used mBART, as previous research showed it outperformed mT5 [10]. Translation-specific models such as NLLB [21] were omitted, because they cannot be further pre-trained.
Languages: We selected the LRL gu, kn, si, ta along with en and hi (see Table 1). kn is not included in mBART and is at a disadvantage compared to other languages [22]. To fine-tune mBART for kn, we used related language fine-tuning strategy [6] with Telugu.
@@ 4 Experimental Setting
Datasets: We used both open-domain (CC) and domain-specific (Bible, Gvt, PMI) corpora (see Table 3). Each language has parallel data from three domains. The rather small FLORES dataset is used for out-domain testing only. Bible is the only domain-specific corpus with data for all our languages. The PMI and Gvt corpora are mutually exclusive in our experiments. Therefore, when describing results, we use PMI/Gvt to denote that we use one of these corpora for the considered experiment. Unlike PMI/Gvt and Bible, CC is larger, even for LRLs. However, it is a web-mined corpus, and is considered noisy [24, 25]. For multi-domain experiments, we randomly shuffled the training data from different domains.
Evaluation Metrics: We used SentencePiece BLEU (spBLEU) that was used in FLORES-101 evaluation benchmark [26].
Domain Divergence: We used Jenson-Shanon Divergence (JSD) to quantify domain divergence. Kashyap et al. [27] showed that JSD is a very reliable measure to analyze the performance of a model in a new domain. JSD values between our parallel corpora are given in Table 2.
@@
Test set specifications are in Table 3. Further details on the data sets, model training, evaluation metrics (including spBLEU signature), and JSD calculation are given in A. We carried out experiments across techniques, languages, and domains for all En-XX pairs. XX-En experiments were carried out for si and kn only to keep the number of experiments at a manageable level.
@@
5 Results and Discussion
5.1 Pre-training with Bi-text Denoising
Three languages were selected for experiments, considering their representation in mBART (Table 1): kn-not included in mBART, si- 243M tokens and hi- 1757M tokens. We tested for data sizes 1k, 10k, and 25k from individual domains and report the average results. Table 4333see Appendix B for raw results shows that neither of the pre-training techniques was able to surpass the results of vanilla FT on mBART, except in two cases. Even in those two cases, continually pre-trained mBART shows only marginal gains over original mBART.
Interestingly, additional pre-training performed worst for Kannada, which is unseen during the original pretraining of mBART. These results imply that bitext denoising is not a suitable strategy for utilizing auxiliary domain parallel data when the data size is small. Reid and Artetxe [11] reported results might improve if additional monolingual data is incorporated in large quantities; however, this is outside the scope of our research and we leave this for future research. Since pre-training only with parallel data did not yield satisfactory results for vanilla FT, we did not carry out further experiments combining pre-training and ITTL.
5.2 Impact of Data Set Size and Domain Divergence on FT and ITTL
In these experiments, we varied the data set size of both the intermediate and final stages. The smallest size is 1k (termed “small”) and the maximum size is 25k (termed “large”)444These terms for dataset sizes are simply relative. In Ranathunga et al.’s [28], even 25k parallel data is considered an extremely small data setup. However, the maximum data size we can use is bound by domain data availability., where 25k is the largest common size across all three domains (see Table 3). Note that we reused the experiments results of single ITTL of our previous research [10]. For clarity, we used a selected set of EN-XX experiment results for our discussion; full results for EN-XX and XX-EN are in Appendix C.
To identify the best-performing technique under different data setups, we analyzed the results with respect to dataset sizes of the intermediate (IM) and final (FI) stages in the following four combinations (recall that Adelani et al. [8] considered only the large-small data setup): 1) small-small (IM data 1k - TD data 1k); 2) large-small (IM data 25k - TD data 1k); 3) small-large (IM data 1k - TD data 25k); 4) large-large (IM data 25k - TD data 25k).
5.2.1 In-Domain Test Case

### v2

@@ Abstract.
1. Introduction
@@
The need to rapidly develop translation systems to disseminate information during pandemics, natural disasters, and other major events justifies the need to build domain-specific Neural Machine Translation (NMT) systems with any and all available parallel data. This is particularly true for low-resource languages (LRLs) that do not possess large amounts of parallel data for any single domain (Zhou and Waibel, 2021). Therefore, when building a domain-specific NMT model for a given LRL pair, we face two possible scenarios: the considered (target, aka test) domain does not have any parallel data to be used in training, or it has a little amount of parallel data. In both cases, it is possible to make use of parallel data available for the same language pair, but in different (auxiliary) domains. This is not an unrealistic assumption - according to Tiedemann (Tiedemann, 2012), hundreds of languages have parallel data from more than one domain.
@@
Multilingual sequence-to-sequence Language Models (msLMs) such as mBART (Tang et al., 2021) have been shown to outperform vanilla Transformer models for the LRL-NMT task (Madaan et al., 2020; Thillainathan et al., 2021; Lee et al., 2022). There are two possible ways to exploit auxiliary domain parallel data in building domain-specific LRL-NMT systems on top of msLMs: to further pre-train the msLM with self-supervised training objectives (continuous pre-training) (Reid and Artetxe, 2022) or to fine-tune the msLM with the NMT objective.
@@
As shown in Figure 1, there are different ways to use auxiliary domain data in fine-tuning. These can be categorized as single-stage fine-tuning (FT) and Intermediate Task Transfer Learning (ITTL) (Phang et al., 2018). Similarly, we can introduce another categorization based on the data availability of the target domain. In-domain refers to the case whether target domain has some parallel data for model training. Out-domain refers to the case where there is no such parallel data for the target domain. Note that out-domain testing scenario can be considered as zero-shot testing with respect to the target domain.
@@
Variations of single-stage FT are: 1) single-domain (vanilla) FT, where the msLM is fine-tuned once with data from one domain, and 2) multi-domain FT, where the msLM is fine-tuned once with data from multiple domains. In vanilla FT, if the target domain has parallel data, this data is used for fine-tuning (in-domain case). In multi-domain FT, when the target domain has parallel data (in-domain case), it is combined with auxiliary domain data. In ITTL, the msLM is fine-tuned with an intermediate task, before fine-tuning with the final task. It also has two variations: single-domain ITTL, where the intermediate task has data from one auxiliary domain, and multi-domain ITTL, where the intermediate task has data from multiple domains. In both cases, the final task (i.e. stage111We use the term stage interchangeably with task, i.e. final task or final stage, and intermediate task or intermediate stage.) fine-tuning is done using data from a single domain. If target domain has parallel data (in-domain case), the final task always uses this data. In the in-domain case, target domain data is also combined with auxiliary domain data in the intermediate stage of multi-domain ITTL.
@@
Currently, there is no comparative study on the impact of continuous pre-training and fine-tuning techniques when auxiliary domain data is used. Determining the method that best exploits auxiliary domain parallel data is beneficial for LRL-NMT researchers who are unable to run extensive experiments due to limited computational resources. Lee et al. (Lee et al., 2022) and Adelani et al. (Adelani et al., 2022) identified that domain divergence in parallel datasets affects NMT performance. However, they did not conduct a detailed investigation into the impact of this factor. Khiu et al. (Khiu et al., 2024) and our previous work (Nayak et al., 2023) evaluated the impact of domain divergence only in the case of single-domain ITTL. Given the above shortcomings, this paper intends to answer the following questions related to domain-specific LRL-NMT systems built on top of msLMs:
@@
In order to answer these questions, we conducted an extensive set of experiments with the aforementioned FT techniques as well as the bitext denoising pre-training technique (Reid and Artetxe, 2022). We considered the LRLs Sinhala (si), Tamil (ta), Kannada (ka), and Gujarati (gu) as well as Hindi (hi) for English-centric NMT. We report results for both the in-domain and out-domain cases. We quantify the impact of domain divergence using Jenson-Shanon Divergence (JSD).
@@
When the parallel data set size is small (less than 25k, in our experiments), pre-training with bitext denoising yields no gains for both in- and out-domain setups.
@@
Multi-domain ITTL (with data from two domains in the intermediate step) is the best for in-domain, but its utility diminishes when the target dataset size increases. For out-domain, the best strategy depends on the domain relatedness and the size of parallel corpora.
@@
For both in- and out-domain setups, simply combining data from more than two domains does not help in ITTL, mainly due to the divergence between the domains they belong to.
2. Related Work
2.1. ITTL on msLMs for NMT
@@
ITTL has roots in Transfer Learning, which has been extensively explored for NMT with vanilla Transformer models (Imankulova et al., 2019; Gao et al., 2024). However, as mentioned earlier, for LRLs, NMT systems built on top of msLMs outperform these traditional NMT techniques. Thus we do not discuss those Transfer Learning techniques.
@@
ITTL (Phang et al., 2018; Pruksachatkun et al., 2020; Park and Caragea, 2020) refers to fine-tuning a pre-trained language model with the intermediate task(s), before fine-tuning on the final task. It has been extensively experimented with encoder-based LLMs such as BERT (Phang et al., 2018; Moghe et al., 2021), and has shown promising results even for LRLs (Dhananjaya et al., 2024).
@@
With respect to ITTL for NMT, the closest to our work is Adelani et al. (Adelani et al., 2022). Similar to us, they explored single-domain ITTL and multi-domain ITTL, as well as multi-domain FT. However, they only considered the scenario where the auxiliary task has a large amount of data, while the target task has only a small amount of data. Moreover, they only used religious and news data and tested only for the in-domain setup. They did not quantify the impact of domain divergence. Our prior work (Nayak et al., 2023) and Khiu et al. (Khiu et al., 2024) considered only single-domain ITTL. However, they considered both in- and out-domain test setups and investigated the impact of domain divergence. Verma et al. (Verma et al., 2022) experimented with single-domain ITTL and multilingual ITTL (data from the same domain but in multiple languages). They found that the former is better, but did not quantify the impact of domain divergence.
2.2. Pre-training msLMs with Parallel Data
@@
Reid and Artetxe (Reid and Artetxe, 2022) augmented the existing denoising objective in mBART with three new objectives: replace words in the noised sequence with a bilingual dictionary, predict the reference translation instead of the input sequence (bitext denoising), and a combination of the two. Kale et al. (Kale et al., 2021) used four denoising objective functions on mT5: translation language modeling, standard NMT, denoised NMT and denoised NMT with language model (LM). Chi et al. (Chi et al., 2021) used three cross-lingual objective functions on mT5: machine translation, translation pair span corruption, and translation span corruption. However, the latter two research did not consider the NMT task.
2.3. NMT for Low-resource Languages
In the context of the task of Machine Translation, a low-resource language refers to a language for which there are no sufficient parallel corpora to train Machine Translation models (Ranathunga et al., 2021). Thus, NMT results obtained for high-resource languages such as English and French cannot be obtained for such low-resource languages (Haddow et al., 2022). Therefore, several prominent solutions have been introduced over the past years (Haddow et al., 2022; Ranathunga et al., 2021). As mentioned earlier, Transfer Learning is one such technique. Other techniques include unsupervised (Artetxe et al., 2018) and semi-supervised NMT (Cheng et al., 2016), data augmentation (Tennage et al., 2018) and pivoting (Cheng, 2019).
Early NMT solutions were implemented on either recurrent neural network architectures, or the Transformer architecture (Vaswani et al., 2017), where the model is trained from scratch, for the task of NMT (Imankulova et al., 2019). However, with the introduction of pre-trained Language Models such as BERT (Devlin et al., 2019) and BART (Lewis et al., 2020) (as well as their multilingual counterparts mBERT and mBART), researchers were keen on exploiting such models in the context of low-resource NMT, to benefit from the cross-lingual transfer abilities of such models. While encoder models such as BERT were used as language model priors (Baziotis et al., 2020), encoder-decoder models such as mBART could be directly fine-tuned for the NMT task (Thillainathan et al., 2021). With the revelation that such encoder-decoder models significantly outperform the vanilla Transformer models for low-resource NMT (Lee et al., 2022), encoder-decoder models such as NLLB ( No Language Left Behind) (Costa-jussà et al., 2022), which have been specifically trained for the task of NMT were released. With the introduction of Large Language Models (LLMs) such as ChatGPT (OpenAI, 2023) and Claude (Anthropic, 2024), researchers have conducted empirical studies on their effectiveness against the encoder-decoder models. However, LLMs have not been able to fully outperform encoder-decoder models such as NLLB, in the context of low-resource languages as yet (Enis and Hopkins, 2024; Robinson et al., 2023).
3. Methodology
@@ 3.1. Continuous Pre-training
For continuous pre-training, we experimented with two of the Reid and Artetxe’s (Reid and Artetxe, 2022) pre-training objective functions: bitext de-noising with auxiliary domain parallel data alone (), and combined with monolingual denoising () (Reid and Artetxe, 2022). For a given source-target sentence pair (), , where the likelihood of generating the target sentence is optimized, conditioned on the noised version of the source sentence . The noising function removes spans of text and replaces them with a mask token. For monolingual denoising, we used the source and target side of the parallel data separately and did not use additional monolingual data as done by Reid and Artetxe (Reid and Artetxe, 2022), because our objective is only to exploit the use of parallel data. We also excluded their dictionary denoising, as many LRLs still do not have sufficient bilingual dictionaries (Liyanage et al., 2021). During pre-training, msLM is further trained, starting from its current weights. Once msLM is pre-trained with Reid and Artetxe’s objective functions (Reid and Artetxe, 2022), we continue with vanilla FT using the NMT objective function.
3.2. Fine-tuning
We use the four FT strategies (a, b, c, d) shown in Figure 1, where strategies c and d are extensions of strategies a and b, respectively:
Fine-tune the msLM with data from a single domain
Fine-tune the msLM with data from multiple domains
First fine-tune the msLM with data from a single domain, then fine-tune the resulting model with data from another single domain
First fine-tune the msLM with data from multiple domains, then fine-tune the resulting model with data from a single domain
All parameters of the encoder, decoder, and embedding layers were updated during all the fine-tuning strategies; i.e. no components were frozen. The model retained the original architecture with pre-layer normalization and embedding layer normalization.
@@ 3.2. Fine-tuning
We formally define these strategies as follows: Let be the set of all domains where there is parallel data available for model training for a considered language pair. is a domain for which there is no training data, but only test data is available. In the single-domain FT (vanilla FT) baseline (Figure 1(a)), the model is only fine-tuned with data from one domain . In the in-domain case, both fine-tuning and testing is done using data from domain . In the out-domain case, fine-tuning uses data from , but testing is done on domain . In multi-domain FT (Figure 1(b)), the model is fine-tuned with data from multiple domains . In the in-domain case, this includes data from the target domain as well. In the out-domain case, this trained model is tested on the unseen target domain . In single-domain ITTL (Figure 1(c)), msLM is first fine-tuned with data of domain . This is the intermediate task, which is also called the first stage. Next, this fine-tuned model is further fine-tuned with data from domain . This is the second and last stage222The model can always be fine-tuned in more stages if there is data from different domains. However, we do not consider this option due to lack of computational resources.. In the in-domain case, testing will be done on the same domain . In out-domain case, testing will be done on the unseen domain . In multi-domain ITTL (Figure 1(d)), msLM is first fine-tuned with data from multiple domains in . This model is further fine-tuned with , which is the target domain. In the in-domain case, data of is included in the intermediate stage, as well as in the final stage. In the out-domain case, testing will be done on the unseen domain .
4. Experimental Setup
Our previous research was one of the first to explore the benefits of single-domain ITTL for LRL-NMT, considering the divergence across domains. We reused the experimental results of single-domain ITTL (strategy c in Figure 1) of our previous research (Nayak et al., 2023). We used the same set of languages and the datasets as in our previous research, and since it experimentally showed that for those languages, mBART outperformed mT5 (Nayak et al., 2023), we only experimented with the former. NLLB (Costa-jussà et al., 2022) was used for an ablation study.
We selected the LRL gu, kn, si, ta along with en and hi (see Table 1). kn is not included in mBART and is at a disadvantage compared to other languages (Ranathunga and de Silva, 2022). To fine-tune mBART for kn, we used related language fine-tuning strategy (Lee et al., 2022) with Telugu being the related language333mBART expects the input sentences to be annotated with the language ID. However, these IDs should be from the list of languages that mBART supports. Since Kannada is not supported in mBART, we used the language ID of Telugu to annotate Kannada sentences.. In order to examine whether this language selection has an impact on Kannada results, we carried out an ablation study using Tamil instead of Telugu. As shown in Table 19 in Appendix C, this language selection shows no impact on the Kannada results.
@@ 4. Experimental Setup
Datasets: We used both open-domain (CC) and domain-specific (Bible, Gvt, PMI) corpora (see Table 3). Each language has parallel data from three domains. The rather small FLORES dataset is used for out-domain testing only. Bible is the only domain-specific corpus with data for all our languages. The PMI and Gvt corpora are mutually exclusive in our experiments. Therefore, when describing results, we use PMI/Gvt to denote that we use one of these corpora for the considered experiment. Unlike PMI/Gvt and Bible, CC is larger, even for LRLs. However, it is a web-mined corpus, and is considered noisy (Kreutzer et al., 2022; Ranathunga et al., 2024). For multi-domain training, we randomly shuffled the training data from different domains.
Evaluation Metrics: We used SentencePiece BLEU (spBLEU) that was used in FLORES-101 evaluation benchmark (Goyal et al., 2022).
Domain Divergence: We used Jenson-Shanon Divergence (JSD) to quantify domain divergence. Kashyap et al. (Ramesh Kashyap et al., 2021) showed that JSD is a very reliable measure to analyze the performance of a model in a new domain. JSD values between our parallel corpora are given in Table 2.
@@
Test set specifications are in Table 3. Further details on the data sets, model training, evaluation metrics (including spBLEU signature), and JSD calculation are given in Appendix A. We carried out experiments across techniques, languages, and domains for all En-XX pairs. XX-En experiments were carried out for si and kn only to keep the number of experiments at a manageable level.
@@
5. Results and Discussion
5.1. Pre-training with Bi-text Denoising
Three languages were selected for experiments, considering their representation in mBART (Table 1): kn-not included in mBART, si- 243M tokens and hi- 1757M tokens. We tested for data sizes 1k, 10k, and 25k from individual domains and report the average results. Table 4444see Appendix B for raw results. shows that neither of the pre-training techniques was able to surpass the results of vanilla FT on mBART, except in two cases where only marginal gains are reported over the original mBART.
Interestingly, additional pre-training performed worst for Kannada, which is unseen during the original pretraining of mBART. These results imply that bitext denoising is not a suitable strategy for utilizing auxiliary domain parallel data when the data size is small. Reid and Artetxe (Reid and Artetxe, 2022) reported results might improve if additional monolingual data is incorporated in large quantities; however, this is outside the scope of our research and we leave this for future research. Since pre-training only with parallel data did not yield satisfactory results for vanilla FT, we did not carry out further experiments combining pre-training and ITTL. However, it is possible to further investigate the impact of pre-training, by varying the dataset size, and the noising rate, etc.
5.2. Impact of Data Set Size and Domain Divergence on FT and ITTL
In these experiments, we varied the data set size of both the intermediate and final stages. The smallest size is 1k (termed “small”) and the maximum size is 25k (termed “large”)555These terms for dataset sizes are simply relative. In Ranathunga et al.’s (Ranathunga et al., 2021), even 25k parallel data is considered an extremely small data setup. However, the maximum data size we can use is bound by domain data availability., where 25k is the largest common size across all three domains (see Table 3). For clarity, we used a selected set of EN-XX experiment results for our discussion; full results for EN-XX and XX-EN are in Appendix C.
To identify the best-performing technique under different data setups, we analyzed the results with respect to dataset sizes of the intermediate (IM) and final (FI) stages in the following four combinations (recall that Adelani et al. (Adelani et al., 2022) considered only the large-small data setup): 1) small-small (IM data 1k - FI data 1k); 2) large-small (IM data 25k - FI data 1k); 3) small-large (IM data 1k - FI data 25k); 4) large-large (IM data 25k - FI data 25k).
5.2.1. In-Domain Test Case